OpenStudy (anonymous):

Select the factors of x2 − 14x + 49. (x + 49)(x + 1) (x + 7)(x + 7) (x − 7)(x − 7) (x − 49)(x − 1)

OpenStudy (anonymous):

To find the factors, you want to find two numbers that multiply to 49 and add to -14. Knowing that, it would (x-7)(x-7)

OpenStudy (anonymous):

the answer is C yes

OpenStudy (anonymous):

Looking at the expression x%5E2-14x%2B49, we can see that the first coefficient is 1, the second coefficient is -14, and the last term is 49. Now multiply the first coefficient 1 by the last term 49 to get %281%29%2849%29=49. Now the question is: what two whole numbers multiply to 49 (the previous product) and add to the second coefficient -14? To find these two numbers, we need to list all of the factors of 49 (the previous product). Factors of 49: 1,7,49 -1,-7,-49 Note: list the negative of each factor. This will allow us to find all possible combinations. These factors pair up and multiply to 49. 1*49 = 49 7*7 = 49 (-1)*(-49) = 49 (-7)*(-7) = 49

OpenStudy (anonymous):

post all ur questions here if u wish

OpenStudy (zaarin):

(x-7)(x-7)

OpenStudy (anonymous):

Factor Completely: 25x2 − 36 (25x − 6)(x + 6) (25x + 6)(x − 6) (5x + 6)(5x − 6) (5x − 6)(5x − 6)

OpenStudy (zaarin):

but the answer is the third one

OpenStudy (anonymous):

(5x)^2-6^2 =(5x+6)(5x-6)

OpenStudy (anonymous):

The third one. You can double check by using foil 25x-30x+30x - 36 25x - 36

OpenStudy (anonymous):

Identify the GCF of 9a4b4 − 27a3b3 + 18a3b2. 9a3b3 9a3b2 3a2b3 3a3b2

OpenStudy (anonymous):

thanks so much samsterz <3

OpenStudy (zaarin):

9a3b2

OpenStudy (anonymous):

9a3b2 9a^4b^4 - 27a^3b^3 + 18a^3b² = 9a^3b² ( ab² - 3b + 2 )

OpenStudy (anonymous):

Select the factors of xy − 4x + 2y − 8. (x − 2)(y + 4) (x + 2)(y − 4) (x − 4)(y + 2) (x + 4)(y − 2)

OpenStudy (anonymous):

(x + 2)(y − 4)

OpenStudy (zaarin):

xy+2y-4x-8 y(x+2)-4(x+2) (y-4)(x+2)

OpenStudy (anonymous):

Find the factors of 2a2 + 9a − 18. (2a − 3)(a + 6) (2a + 3)(a − 6) (2a − 9)(a + 2) (2a − 2)(a + 9)

OpenStudy (anonymous):

answer is A

OpenStudy (anonymous):

next question ill go faster now since im concentrate on ur questions fully now

OpenStudy (anonymous):

Determine the factors of a2b + 2a2 + 3b + 6. (b + 3)(a2 + 2) (b + 2)(a2 + 3) (ab + 2)(a + 3) (ab + 3)(a + 2)

OpenStudy (anonymous):

2a^2?

OpenStudy (anonymous):

yes

OpenStudy (zaarin):

a2(b+2)+3(b+2) (a2+3)(b+2)

OpenStudy (anonymous):

looks like hes right

OpenStudy (zaarin):

its a she not a he btw

OpenStudy (anonymous):

thats not a choice .. & lol

OpenStudy (anonymous):

a^2b?

OpenStudy (anonymous):

is how it starts correct?

OpenStudy (anonymous):

ohhh. yeah!

OpenStudy (anonymous):

check lolol

OpenStudy (anonymous):

Identify the factors of x2 − 4x − 12. (x + 4)(x − 3) (x − 4)(x + 3) (x − 2)(x + 6) (x + 2)(x − 6)

OpenStudy (anonymous):

the other one was right :)

OpenStudy (anonymous):

a^2 + 3)(b + 2)

OpenStudy (anonymous):

(x+2)(x-6)

OpenStudy (anonymous):

Factor completely: 2x3 + 6x2 + 2x + 6 2[(x + 3)(x2 + 1)] (2x + 6)(x2 + 1) (x + 3)(2x2 + 2) 2(x3 + 3x2 + x + 3)

OpenStudy (anonymous):

2(x^2+1)(x+3)

OpenStudy (anonymous):

or

OpenStudy (anonymous):

(2x + 6)(x2 + 1)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

lolol what a deceitful bunch of answer choices lol

OpenStudy (anonymous):

Lmao

OpenStudy (anonymous):

next question

OpenStudy (anonymous):

Factor x2 − 8x + 9 (x − 9)(x − 1) (x + 9)(x + 1) Prime (x + 9)(x - 1)

OpenStudy (anonymous):

prime

OpenStudy (anonymous):

Factor completely: −5x2 + 10x − 15 −5(x2 − 2x + 3) −1(x2 − 10x + 15) −5x(x2 − 2x + 3) −5(x2 − 2x − 3)

OpenStudy (anonymous):

-5(x^2-2x+3)

OpenStudy (anonymous):

What are the factors of 3x2 + 10x + 8? (3x + 8)(x + 1) (3x + 1)(x + 8) (3x + 2)(x + 4) (3x + 4)(x + 2)

OpenStudy (anonymous):

(x+2)(3x+4)

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

Which polynomial is a difference of two squares? x2 − 12 x2 + 12 x2 − 36 x2 + 36

OpenStudy (anonymous):

i think this one is c .. but idk for sure

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

looks like ur right

OpenStudy (anonymous):

yay i got one right :)

OpenStudy (anonymous):

What are the factors of x2 + 10x + 24? (x − 12)(x − 2) (x + 12)(x + 2) (x + 4)(x + 6) (x − 4)(x − 6)

OpenStudy (anonymous):

(x+6)(x+4)

OpenStudy (anonymous):

Identify the factors of 9x2 + 24x + 16. (3x + 4)(3x + 4) (3x − 4)(3x − 4) (3x + 2)(3x + 8) (3x − 8)(3x − 2)

OpenStudy (anonymous):

(3x+4)(3x+4) but let me check

OpenStudy (anonymous):

yup im right

OpenStudy (anonymous):

Factor completely: 9x4y − 6x3y2 + 3x2y3 3x2y(3x2 − 2xy + y2) 3xy(3x2 − 2xy + y2) 3xy2(3x2 − 2xy + y2) 3x2y(3x2 + 2xy + y2)

OpenStudy (anonymous):

9x^4 space y?

OpenStudy (anonymous):

i only have one question left :D

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

9x 4y right no epxonents?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

write out the equation with the exponents lol

OpenStudy (anonymous):

3x2y(3x2−2xy+y2)

OpenStudy (anonymous):

100% right

OpenStudy (anonymous):

9x^4y − 6x^3y^2 + 3x^2y^3

OpenStudy (anonymous):

o

OpenStudy (anonymous):

What are the factors of x2 − 81? (x + 9)(x − 9) (x − 9)(x − 9) (x + 3)(x − 27) Prime

OpenStudy (anonymous):

last one :))

OpenStudy (anonymous):

(x-9)(x+9)

OpenStudy (anonymous):

What are the factors of x2 − 81? (x + 9)(x − 9) (x − 9)(x − 9) (x + 3)(x − 27) Prime

OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 3x2 − 3x − 60.

OpenStudy (anonymous):

3x2−3x−60 Factor out the GCF of 3 from each term in the polynomial. Factor out the GCF of 3 from the expression 3x2. 3(x2)−3x−60 Factor out the GCF of 3 from the expression −3x. 3(x2)+3(−x)−60 Factor out the GCF of 3 from the expression −60. 3(x2)+3(−x)+3(−20) 3(x2)+3(−x)+3(−20) Factor out the GCF of 3 from 3x2−3x−60. 3(x2−x−20) In this problem 4⋅−5=−20 and 4−5=−1, so insert 4 as the right hand term of one factor and −5 as the right-hand term of the other factor. 3(x+4)(x−5)

OpenStudy (anonymous):

or

OpenStudy (anonymous):

wow you are fantastic at this you should be a teacher no joke lol

OpenStudy (anonymous):

facter out 3 out of all of it.. all you do is you look at the problem, "3x^2-3x-60" looking at that there is a leading coefficient of 3 (3x^2) so you want to get the x^2 by itself, to do that you divide it all by 3 which gives you the 3(x^2-x-20)

OpenStudy (anonymous):

u can also use that if u wish and ur welcome

OpenStudy (anonymous):

what math class r u in?

OpenStudy (anonymous):

i used the other one :) & algebra on virtual school

OpenStudy (anonymous):

o

OpenStudy (anonymous):

gl

OpenStudy (anonymous):

lol thanks! :)

OpenStudy (anonymous):

quadratic equations is what im on now .. -__-

OpenStudy (anonymous):

go ahead post them

OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 8x − 4. (−2, 20) (−2, −12) (2, 20) (−2, −28)

OpenStudy (anonymous):

what grade are you in? :o

OpenStudy (anonymous):

like 11th grade

OpenStudy (anonymous):

im in 10th

OpenStudy (anonymous):

vertex (-2,-12)

OpenStudy (anonymous):

A quadratic equation has an x-squared term. Sometimes Always Never

OpenStudy (anonymous):

sometimes ?

OpenStudy (anonymous):

always

OpenStudy (anonymous):

remeber that formula for quadratic formula and whys its different from exponent adn lineqr formulas

OpenStudy (anonymous):

its becasue it holds an exponent